Substitute natural gas which can be used as a fuel is prepared via the following reaction:
4 CO(g) + 8 H2 (g) 3 CH4 (g) + CO2 (g) + 2H2O(l)

Using the data given below determine ΔHo for the substitute natural gas reaction above.
C (graphite) + ½ O2(g)---------- CO(g) -110.5 kJ
CO(g) + ½ O2(g)------- CO2(g) -283.0 kJ
H2(g) + ½ O2(g)----------- H2O(l) -285.8 kJ
C (graphite) + 2 H2(g)----------CH4(g) -74.81 kJ
CH4(g) + 2 O2(g)---------- CO2(g) + 2 H2O(l) -890.3 kJ
1
Expert's answer
2012-10-03T09:04:40-0400
3CO + 1.5 O2 + 6H2& + 3 O2 + C + 2H2+ CO + 2CO2 + 4H2O----> 3CO2 + 6H2O + 1CH4+1C& +& 1/2O2 +2CH4 +4O2
&
3 X CO(g) + ½ O2(g)------- CO2(g) 3*& -283.0 kJ& = - 849 kJ
6 X H2(g) + ½ O2(g)----------- H2O(l) 6* -285.8 kJ& = - 1714.8 kJ
1 X& C (graphite) + 2 H2(g)----------CH4(g) & -74.81 kJ
1 X CO(g)----------C (graphite) + ½ O2(g) & 110.5& kJ
2 X& CO2(g) + 2 H2O(l)-------------- CH4(g) + 2 O2(g) 2*890.3 = 1780.6 kJ

ΔHo=-849+ (-1714.8) +(-74.81) + 110.5 +1780.6 = -747.51 kJ